OpenStudy (afloridagirl):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0

OpenStudy (afloridagirl):

im not sure what this is even asking..

OpenStudy (afloridagirl):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

have you learned about the term discriminant? http://www.regentsprep.org/regents/math/algtrig/ate3/discriminant.htm

OpenStudy (afloridagirl):

a little bit

OpenStudy (afloridagirl):

b2 − 4ac

OpenStudy (afloridagirl):

i think it was that one but i wanna make sure..

jimthompson5910 (jim_thompson5910):

on that page, it says `If the discriminant is a perfect square, the two roots are rational numbers.` which is another way of saying "if the discriminant is a perfect square, then we can factor the quadratic in terms of whole numbers"

jimthompson5910 (jim_thompson5910):

In the case of 4x^2 + 6x + 2 we have a = 4 b = 6 c = 2 plug those values into b^2 - 4ac and tell me what result you get.

OpenStudy (afloridagirl):

4x^2 + 6x + 2 = 0 4x^2 + 4x + 2x + 2 = 0 4x(x + 1) + 2(x + 1) = 0 (4x + 2)(x + 1) = 0

OpenStudy (afloridagirl):

is that right?

OpenStudy (afloridagirl):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

They don't want you to factor. They just want the value of the discriminant.

OpenStudy (afloridagirl):

???

jimthompson5910 (jim_thompson5910):

`Use the part of the quadratic formula that you chose above and find its value,` they want the value of \(\Large \color{red}{b^2 - 4ac}\)

jimthompson5910 (jim_thompson5910):

a = 4 b = 6 c = 2 \[\Large b^2 - 4ac = 6^2-4*4*2 = ???\]

OpenStudy (afloridagirl):

4?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

extra info: since the discriminant (4) is a perfect square, this means we can factor 4x^2+6x+2 over the whole numbers. You've just shown that above. You didn't need to factor, but it's good practice to factor so it doesn't hurt.

jimthompson5910 (jim_thompson5910):

So as you can guess, the discriminant helps save you a lot of time if you're given something 20x^2 + 15x - 29 which isn't clear if it can be factored or not.

OpenStudy (afloridagirl):

so 4 is the answer?

jimthompson5910 (jim_thompson5910):

There are 2 answers they want here. The first part would be b^2-4ac. The second part would be 4.

jimthompson5910 (jim_thompson5910):

The two are very similar and linked together

OpenStudy (afloridagirl):

so how do i put that?

jimthompson5910 (jim_thompson5910):

What do you mean?

OpenStudy (afloridagirl):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0 Numerical Answers Expected!

OpenStudy (afloridagirl):

i have to type it in

jimthompson5910 (jim_thompson5910):

So it sounds like they just want you to type in a single number

jimthompson5910 (jim_thompson5910):

based on the statement `Numerical Answers Expected!`

jimthompson5910 (jim_thompson5910):

I thought it wanted you to pick from \(\Large -b\) or \(\Large b^2 - 4ac\) or \(\Large 2a\) but I guess they don't want that as part of the answer.

OpenStudy (afloridagirl):

thats why i said i didnt know what to do

jimthompson5910 (jim_thompson5910):

if it were me, I'd just type in the single number result you got

jimthompson5910 (jim_thompson5910):

it sounds like that's all they want to know

OpenStudy (afloridagirl):

okay i will try that